OpenStudy (anonymous):

Write the trigonometric expression in terms of sine and cosine, and then simplify. tan(x)csc(x)

OpenStudy (anonymous):

tan(x)=sin(x)/cos(x) csc(x)=1/sin(x) tan(x)csc(x)=1/cos(x)